Two capacitors with 2.0 F and 3.0 F capacitance, respectively, are connected in series and subjected to a total potential difference of 100 V.



Find the


(a) total capacitance,



(b) charge stored in each capacitor, and



(c) potential difference across each capacitor.

1
Expert's answer
2022-03-24T11:00:58-0400

(a) total capacitance

"C=\frac{C_1C_2}{C_1+C_2}"

"C=\frac{2.0*3.0}{2.0+3.0}=1.2\:\rm F"

(b) charge stored in each capacitor

"q=CV\\
=1.2*100=120\:\rm C"

(c) potential difference across each capacitor

"V_1=q/C_1=120/2.0=60\:\rm V"

"V_2=q/C_2=120/3.0=40\:\rm V"